OpenStudy (anonymous):

i have a question can someone will help me with my calculus and vectors assignment

OpenStudy (amistre64):

if you post the actual information we might be able to determine if we can or cannot ....

OpenStudy (anonymous):

find f^(3) (x) if f(x) = 1/sqrt(3X + 4)

OpenStudy (amistre64):

the third derivative of that function right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

id rewrite it as:\[(3x+4)^{-1/2}\] and work it from that angle

OpenStudy (anonymous):

but how i can solve it

OpenStudy (amistre64):

the derivative rules are pretty much basic and self explanatory. If you do not yet know them then it might be good for you to revisit them.

OpenStudy (amistre64):

\[\frac{d}{dx}x^n=nx^{n-1}\] \[\frac{d}{dx}f(g(x))=f'(g(x))*g'(x)\] \[\frac{d}{dx}fg=f'g+fg'\] those look like the 3 that you might want to apply to this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

find the equations of the tangents to the curves f(x) = x^2/3 and g(x)=9/x at their point of intersection.

OpenStudy (amistre64):

an equation of a tangent line needs a point and a slope to be defined. since slope IS derivative we would need to determine the derivative of each function.

OpenStudy (amistre64):

then we would need to determine the point at which they actually intersect to define the "point"

OpenStudy (amistre64):

y = f'(xo)x -f'(xo)xo + yo y = g'(xo)x -g'(xo)xo + yo given the point of intersection as (xo,yo)

OpenStudy (anonymous):

a ball is thrown upward and forward so that after time t (in seconds) its height is h(t) = 20t-5t^2. After starting to descent it strikes a wall at a point 15m high. find the rate of descent or downward velocity of the ball at the moment of impact against the wall.

OpenStudy (amistre64):

new questions should be posted on the left in order to allow more people to have an opportunity to respond to them.

OpenStudy (anonymous):

oh...sorry...i didn't know that